OpenStudy (anonymous):

Find the values of x and y: x+y+xy=-5 x^2+y^2+x^2y^2=49

OpenStudy (anonymous):

Can someone please help me?

hartnn (hartnn):

yes x+y = -5 - xy square both sides, what do u get ?

OpenStudy (anonymous):

x^2+y^2=25-x^2y^2

hartnn (hartnn):

\((a+b)^2 = a^2+2ab+b^2 \)

hartnn (hartnn):

so, (x+y)^2 =..?

OpenStudy (anonymous):

x^2+y^2+2xy

hartnn (hartnn):

what about (-5-xy)^2 ?

OpenStudy (anonymous):

25+x^2y^2+10xy

hartnn (hartnn):

yes, so you get x^2+2xy+y^2 = 25 + x^2y^2 + 10xy right ?

OpenStudy (anonymous):

Yes

hartnn (hartnn):

now from 2nd equation x^2+y^2+x^2y^2=49 we get the value of x^2+y^2 as 49 - x^2y^2

hartnn (hartnn):

so, x^2+2xy+y^2 = 25 + x^2y^2 + 10xy changes to 49 -x^2y^2 + 2xy = 25 +x^2y^2 +10xy this equation is a Quadratic in 'xy' like if you put p=xy 49 - p^2 +2p = 25 + p^2 +10p can you solve this quadratic in 'p' ? to get 2 values of p (which are actually 2 values of xy)

OpenStudy (anonymous):

ooooohhhhh

hartnn (hartnn):

let me give you steps , and see whether you get it 49 -25 +2p - 10 p - p^2-p^2 = 0 so, 24 - 8p -2p^2 = 0 which is p^2 +4p -12 = 0

hartnn (hartnn):

that quatic can be factored easily to get 2 values of p p^2 + 6p - 2p -12 = 0 p(p+6) - 2 (p+6) = 0 (p-2)(p+6) = 0 p=2, -6 so the 2 values of xy are 2 and -6

hartnn (hartnn):

**quadratic

hartnn (hartnn):

so, x^2y^2 has 2 values as +4 and +36

hartnn (hartnn):

using this : x+y = -5 - xy we can get 2 values of x+y too

hartnn (hartnn):

and if we know x+y and xy, we can get x and y easily try it, if you don't get i'll help later (then you also ask if you have doubts in the above explanation)

OpenStudy (anonymous):

Umm thanks. I'll try it again.

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

Hello

hartnn (hartnn):

could you get the values ? which step are you stuck ?